What is the expected value?

Outcome

Value of

Random Variable

Probability

A

1

.4

B

2

.3

C

3

.2

D

4

.1

A.) 2.5

B.) 10

C.) 3

D.) 0.75
Answer #1

From given table,

Mean = X * P(X)

= 1 * 0.4 + 2 * 0.3 + 3 * 0.2 + 4 * 0.1

= 0.4 + 0.6 + 0.6 + 0.4

= 2

Expected value = 2

Not in the option value of 2